### v3

# What-if Analysis for Business Users: Current Practices and Future Opportunities

Abstract.
What-if analysis (WIA), crucial for making data-driven decisions, enables users to understand how changes in variables impact outcomes and explore alternative scenarios. However, existing WIA research focuses on supporting the workflows of data scientists or analysts, largely overlooking significant non-technical users, like business users. We conduct a two-part user study with 22 business users (marketing, sales, product, and operations managers). The first study examines existing WIA techniques employed, tools used, and challenges faced. Findings reveal that business users perform many WIA techniques independently using rudimentary tools due to various constraints. We implement representative WIA techniques identified previously in a visual analytics prototype to use as a probe to conduct a follow-up study evaluating business users’ practical use of the techniques. These techniques improve decision-making efficiency and confidence while highlighting the need for better support in data preparation, risk assessment, and domain knowledge integration. Finally, we offer design recommendations to enhance future business analytics systems.
Business Intelligence, What-if Analysis, Predictive and Prescriptive Analytics, Interview Study
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Conference on Human Factors in Computing Systems; April 26–May 01, 2025; Yokohama, Japanisbn: 978-1-4503-XXXX-X/18/06submissionid: 2546ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Visual analyticsccs: Human-centered computing Graphical user interfaces
1. Introduction
What-if analysis (WIA) is crucial for making data-driven decisions (Gathani et al., 2022; Dimara et al., 2021). For example, business managers may want to understand which marketing activities (e.g., sending emails, making phone calls, giving out free trials, running ad campaigns) should be prioritized to increase customer acquisition rates (Gathani et al., 2022); public health researchers might be interested in the effect of changes in socio-economic factors (e.g., smartphone use is maximized or sleep quality is low) on adolescent obesity rates (Kim et al., 2019); operations analysts may want to identify the optimal time complexities of various sub-algorithms used in a solver (e.g., cut generation, global constraint, local search move) needed to reduce the solver’s overall computation time (Van Cauwelaert et al., 2017). To answer these questions, WIA is essential for identifying key data variables that influence specific outcomes, simulating various scenarios to understand how changes in data variables impact predictions of outcomes, and determining optimal data variable values needed to achieve desired target outcomes. For example, in marketing, given a target outcome of increasing customer acquisition by 5%, WIA can be accomplished by first identifying key variables such as running ad campaigns and making phone calls, building models that predict customer acquisitions based on these variables, and trying alternative scenarios involving different numbers of campaigns and phone calls. Informed decisions can then be made based on the observations on the resulting acquisition rates in these scenarios.
In such examples across many different domains, WIA plays an important role that cannot be fulfilled by other data analysis methods. Similar to many analytic methods, WIA involves building models (e.g., statistical, probabilistic, neural network, optimization, etc.) of independent and dependent variables. However, unlike confirmatory methods such as hypothesis testing, WIA is more exploratory in nature: it involves changing variable values in multiple scenarios and comparing possible outcomes without significance testing. On the other hand, exploratory data analysis (EDA) (Tukey, 1977), which is concerned with understanding trends and patterns in a dataset, is not appropriate for simulating alternative scenarios and predicting their outcomes (Dimara et al., 2021; Oral et al., 2023) to make data-driven decisions. With data growing exponentially (Duarte, 2024), limitations of human memory and cognitive capacity in generating hypotheses, testing scenarios, and predicting outcomes also exacerbate (Gathani et al., 2022). WIA thus addresses an important need that is not supported by methods described, and a growing body of work is dedicated to developing WIA tools for various domains (Wexler et al., 2019; Grafberger et al., 2023; Hohman et al., 2019; Sacha et al., 2018; Gallo et al., 2018; Torsney-Weir et al., 2018; Eric Newburger, 2023).
Despite these efforts to support WIA, existing research mostly focuses on technical users like data analysts and scientists, and overlooks non-technical users who typically lack the expertise or background in coding, statistics, and algorithmic modeling like healthcare experts, urban planners, or law officials (Gathani et al., 2022; Crisan and Correll, 2021; Dimara et al., 2021; Oral et al., 2023). WIA is often practiced by non-technical users because they possess critical domain knowledge that technical users may lack. For example, in WIA, it is often important to brainstorm alternative scenarios and account for potential trade-offs between variables (Dimara et al., 2021), which require extensive domain-specific experiences and intuitions that are not found in professional data analysts. Additionally, non-technical users may also build upon and extend data-driven insights generated by technical users towards specific desired outcomes (Crisan and Correll, 2021; Bhattacharya et al., 2024). Understanding their current WIA practices and pain points is essential for developing better tools that effectively integrate domain knowledge with WIA techniques.
In this paper, we choose to focus on a specific key group of non-technical users conducting WIA: business users [^1] In the remainder of the paper, we use the terms business users and users interchangeably.—such as sales managers, marketing managers, product managers, and operations managers. These users represent a significantly large and influential demographic group (Report, 2022b, a) that regularly makes data-driven decisions without formal training in technical fields like computer science, machine learning, or statistics. They are also the primary audience for self-service business intelligence (BI) tools which constitute a multi-billion dollar industry (Insights, 2024) and stand as the most prominent commercial application of visual data analytics (Software, 2024). Since little is known about how business users employ WIA and the challenges they encounter, we seek to answer the following research questions:
RQ1: What techniques and tools do business users employ to perform WIA for making data-driven decisions?
RQ2: How do they perceive and interact with advanced WIA techniques, and what future opportunities exist for improving these techniques?
We answer these by conducting a two-part user study with 22 business users. For RQ1, we conduct a semi-structured interview study to understand their WIA workflows and challenges when making data-driven decisions. In addition to discussing their own experiences, we ask participants to also walk us through a common business use case of maximizing sales by making spending decisions in various advertising channels, often referred to as marketing mix modeling. Our findings reveal that business users employ various WIA techniques, but rely on rudimentary methods like spreadsheets which are insufficient for making data-driven decisions. Compared to prior studies on data workers’ generic analysis practices (Dimara et al., 2021, 2021; Tory et al., 2021), our study uncovers a specific set of WIA techniques adopted by business users (e.g., sensitivity analysis, driver importance analysis, segmentation analysis) and the associated challenges they face. Moreover, business users prefer performing WIA independently, without relying on data analysts, due to (1) the limited availability of data analysts in enterprises, (2) the inefficiency of communicating with data analysts, (3) the business pressure to make quick decisions, and (4) the importance of incorporating domain knowledge, which data analysts often lack. These findings suggest that currently there is inadequate support for business users to perform WIA making data-informed decisions.
The first interview study informs us of various WIA techniques employed by business users and the challenges they encounter with them. We address RQ2 by first implementing four representative WIA techniques identified in the interview study (i.e., driver importance analysis, sensitivity analysis, goal-seeking analysis, and constrained analysis) in a visual analytics prototype. We then conduct a follow-up study with the same participants, asking them to use this prototype as a probe to make decisions for the same marketing mix modeling use case as in the interview study. Through the study, we identified the potential benefits and opportunities of WIA techniques for business users. Our findings show that they improve business users’ decision-making speed and confidence, highlighting their strong interest in adopting additional advanced analytics techniques for informing their decisions. However, participants also expressed opportunities for more support in preparing data (e.g., consolidating data from multiple sources, consistently defining KPI goals), assessing prediction risks (e.g., understanding confidence and trust in the predictions), and incorporating domain knowledge (e.g., capturing limited budgets, volatile ecological and market conditions such as COVID, war, and inflation).
In summary, this paper contributes:
  • An interview study identifying WIA techniques, tools, and challenges that business users face when making data-driven decisions.
  • A task-based study gathering business users’ hands-on experience and feedback on the benefits and opportunities of using and enhancing WIA techniques for their use in making data-driven decisions.
  • Design recommendations for future business analytics systems to better support business users in advanced analytics.
2. Related Work
Our work relates to prior research on other users performing data analysis, use of enterprise dashboards, advanced analytics tools, and organizational decision-making.
2.1. Understanding Users Performing Data Analysis
Prior work has examined how different types of users perform data analysis. We categorize them along two critical dimensions: task (data understanding and sense-making vs. improving outcomes and decision-making) and technical expertise (having formal technical training vs. no formal training), as depicted in Figure 1.
Figure 1. Layout of data-analyzing users studied in the literature. We map the user groups considering the dimensions of task (x-axis) and technical expertise (y-axis). Our research focuses on business users, who are domain experts lacking technical (e.g., programming, statistics, data science, etc.) background but using data analysis to make better business decisions.
Most studies focus on users with technical expertise (quadrant I), like data analysts, data scientists, and business analysts, who mainly support decisions rather than make them. For example, Kandel et al. (Kandel et al., 2012) identify challenges in analysts’ data pipelines, while Alspaugh et al. (Alspaugh et al., 2019) explore challenges in exploratory analysis. Demiralp et al. (Demiralp et al., 2017) study data scientists’ exploratory analysis, and Zhang et al. investigate business analysts’ needs in generating reports (Zhang et al., 2020) and narratives (Zhang et al., 2022). Although we find business users face similar challenges (e.g., the need to bridge the programming gap for analysts who are less savvy in programming, and the challenges of data integration), our findings highlight their need for data-driven WIA to explore alternative scenarios and the unique challenges related to making decisions in terms of time pressure, uncertainty, and domain knowledge. Newburger and Elmquist study statisticians’ decision-making with a focus on visualization (Eric Newburger, 2023). While their work shares some similarities with ours, we delve deeper into the mechanisms, tools, and challenges of their advanced analyses.
Bartram et al. (Bartram et al., 2021) study “data workers” (quadrants III and IV) with less formal expertise who perform data analysis as part of their daily tasks. Previous work often focuses on data workers in non-business domains, such as political science (South et al., 2020), community leadership (Jasim et al., 2021), and law enforcement (Kang and Stasko, 2012), but their tasks are more about data exploration and sense-making (Klein et al., 2006; Russell et al., 1993; Klein et al., 2007; Pirolli and Card, 2005) than making decisions involving simulating alternative scenarios and action selection (Lunenburg, 2010; Bruch and Feinberg, 2017; Chen, 2008; Oral et al., 2023). Current research also often prioritizes unstructured data, like text. For example, Edge et al. (Edge et al., 2018) report on bringing AI within unstructured sources like social media, news, and cyber intelligence. However, a large number of use cases primarily involve structured data tables, which are underrepresented in studies. Bhattacharya et al. (Bhattacharya et al., 2024) comes closest in empowering non-technical domain experts, particularly healthcare professionals, by enhancing AI models through diverse model explanations (local and global). However, their focus is on improving models and predictions rather than supporting decision-making processes. On contrary, this paper examines how business users explore data scenarios using WIA techniques to practically make decisions.
2.2. Use of Dashboards in Enterprises
Previous work investigates the role of visualization and dashboards in enterprise data analysis and decision-making. For example, Newburger and Elmqvist (Eric Newburger, 2023) explore visualization for statisticians, while studies find that business users, like sales managers, frequently use dashboards (Velcu-Laitinen and Yigitbasioglu, 2012; Bera, 2016; Noonpakdee et al., 2018). However, dashboards are mainly used for communicating findings to stakeholders, not for making data-driven decisions. Dimara and Stasko (Dimara and Stasko, 2021) note the lack of studies on decision-making in visualization research, and other works (Dimara et al., 2021) show how visualizations can be integrated into long-term decisions. Yet, business users face challenges in creating dashboards and conducting analysis necessary for making decisions (Sarikaya et al., 2018; Tory et al., 2021) due to incomplete information shown in dashboard and limited interactivity, which prevents users from drilling into the data to explore different scenarios or conduct simulations. Therefore, visualization alone without data-driven analytics is insufficient, leading us to focus on how advanced analytics can support or hinder making decisions.
2.3. Advanced Analytics Tools
Marx et al. (Marx et al., 2012) and Bergeron et al. (Bergeron et al., 1995) studied legacy information systems in organizations or Executive Information Systems (EIS), identifying gaps in decision-making support, especially in drill-down, scenario analysis, and optimization. Van et al. (Van Bruggen et al., 2010) explored systems for marketing management, echoing these limitations. More recent work by Crisan and Correll (Crisan and Correll, 2021) and Bartram et al. (Bartram et al., 2021) also noted these shortcomings, though focused less on business decisions. Zhang et al. (Zhang et al., 2020) and Honeycutt et al. (Honeycutt et al., 2020) emphasized integrating domain knowledge with AI to enhance user trust in intelligent systems. Bhattacharya et al. (Bhattacharya et al., 2024) empowers healthcare workers to utilise AI model explanations to improve model predictions. Oral et al. (Oral et al., 2023) highlight the lack of decision support features like visibility of alternatives, input processing, change model parameters, etc. in existing tools. Our study expands on exploring model-based WIA techniques for making data-driven decisions, highlighting challenges and opportunities for integrating in them domain-specific expertise.
Most human-in-the-loop machine learning research centers on data scientists and analysts, highlighting their openness to automating decisions (Wang et al., 2019) or collaborating with models (Poursabzi-Sangdeh et al., 2021). Crisan and Fiore-Gartland (Crisan and Fiore-Gartland, 2021) discuss automation’s role in routine tasks, rapid prototyping, and democratizing data science, but for data scientists. Our study differs by focusing on how business users with no formal technical backgrounds (hence referred to as non-technical users) perceive advanced WIA, a form of AutoML used for data-driven decisions. Further, we go beyond interviews, collecting empirical data from users actively engaging in hands-on decision-making with these techniques.
Many business intelligence (BI) tools (Ferrari and Russo, 2016; Tableau, 2021b, a; SAS, 2021) and spreadsheet applications offer some advanced analytics, but they primarily focus on descriptive and exploratory tasks, answering the “what” and “why” in data. But, they often lack predictive and prescriptive features, leaving the “now what?” unanswered to plan subsequent actions. Tools like Excel’s SOLVER (Microsoft, 2021a) and GOAL SEEK (Microsoft, 2021b) provide optimization features for desired outputs (Bartram et al., 2021), but manual formula creation remains challenging for business users, which interactive advanced analytics tools aim to address.
2.4. Organizational Decision-Making
Management and organizational theory categorize decisions into three tiers: strategic (executive level), tactical (middle management), and operational (front-line employees) (Ballou, 1992). This work focuses on tactical decisions. Simon’s three-step model (Simon, 1960) (also observed in (Dimara et al., 2021) and used in (Oral et al., 2023))—intelligence (information gathered and problems identified), design (possible solutions for problems generated and evaluated), and choice (best alternative selected)—describes the decision-making process, as does Berisha-Shaqiri’s eight-step model (Berisha-Shaqiri, 2014)-identification of decision’s purpose, information gathering, principles for judging the alternatives, brainstorm and analyze choices, evaluating alternatives, decision execution, and results evaluation. On similar lines, Jun (Jun, 2022) centers on hypothesis formalization through statistical analyses by data scientists. We observe similar steps in our study participants’ decision-making, but we focus deeper into the data they use, the advanced analytics they apply, and the challenges they face in selecting the best alternatives. Perkins et al. (Perkins and Rao, 1990) examine how marketing managers’ experience influences their decisions and the information they use, while Little (Little, 1979) reports on processes followed by marketing managers from over 40 years ago, but with the aid of descriptive and exploratory tools. In all these works, the role of data and analysis is largely simplified or overlooked. In contrast, our study focuses on how data and advanced analytics are employed by managers across varied departments.
3. Overview of the Two Studies
This section outlines the methods and procedures for our two-part study.
Participant ID Company Sector Company Size Participant ID Company Sector Company Size P1 Education 10,001+ P12 SaaS 201-1000 P2 Consumer Goods 1001-5000 P13 Human Resources 51-200 P3 Non-Profit 51-200 P14 Healthcare 1001-5000 P4 Telecommunications 201-1000 P15 Consumer Goods 10,001+ P5 E-learning 201-1000 P16 Food 10,001+ P6 SaaS 10,001+ P17 Logistics, Supply Chain 11-50 P7 Health, Wellness, Fitness 5001-10,000 P18 IT 201-1000 P8 Telecommunications 10,001+ P19 Hospitality 10,001+ P9 Construction 5001-10,000 P20 Healthcare 10,001+ P10 Consumer Goods 201-1000 P21 SaaS 10,001+ P11 Electrical/Electronic Manufacturing 5001-10,000 P22 SaaS 201-1000
Table 1. Study participants hailed from four business departments of companies with diverse sizes and sectoral focuses. Colors in the table refer to departments in which the participants worked: marketing, sales, product, and operations.
Figure 2. Overview of our two-part study. We conduct the study with 22 business users (12 female, 10 male) from four departments: marketing (7), sales (5), product (5), and operations (5). The interview study aims to understand various what-if analysis (WIA) techniques participants employ, the tools they use, and the challenges they face to use them for their data-driven decision-making of both their own business use cases as well as for a specific business use case of Marketing Mix Modeling (MMM). The task-based study aims to understand the same participant’s perception of practically using representative WIA techniques identified in the first study and implemented as a probe within a visual data analytics prototype. Colors refer to the departments in which the participant worked (marketing, sales, product, operations).
3.1. Participants
We recruited 22 business users through an online survey (Anonymous, 2024b) through the User Interviews (Interviewss, 2021) website and company’s mailing lists. Our survey ensured that participants were selected based on three criteria: (1) use organizational, customer, or product data to guide business decisions, (2) make decisions at least weekly, and (3) have no formal training in data science or programming. Final participant pool comprised business users from four departments: marketing (7), sales (5), product (5), and operations (5) managers, from varied company sizes, sectors, and expertise levels, as shown in Table 1.
3.2. Protocol
Each participant completed both parts of the study over 120–140 minutes in two sessions, held on separate days within one to two weeks, as illustrated in Figure 2. All interviews were conducted and recorded online for analysis. Participants received a total of \$175 in Amazon gift card for completing both sessions. Each study is detailed in later sections.
3.3. Analysis
We analyzed the recorded and transcribed data from both parts of the study using iterative coding to derive our findings. In the first interview study, a researcher initially categorized participants’ business decisions, data usage, and general WIA techniques and tools. The coding then focused on specific WIA techniques, such as scenario exploration and ranking correlations between drivers and key performance indicators (KPIs) among many others, to understand their usage and challenges. Finally, both these analyses were then synthesized to form the overall findings of the interview study.
In the follow-up task-based study, participants’ experiences of practically using four WIA techniques identified from the previous study and implemented within a visual analytics prototype as a probe are captured. Their feedback on potential benefits and improvement opportunities of WIA techniques were similarly categorized. As new data emerged, these categories were refined, with any ambiguities addressed through team discussions. We support our findings with representative participant quotes throughout the paper.
4. Study I: Interview Study
We present the method and findings of the first interview study.
4.1. Method
The interview study was conducted through semi-structured interviews with business users to understand various WIA techniques they employed, motivations behind their use, tools they relied on, and challenges they encountered. First, we describe our method as outlined in Figure 2.
4.1.1. Protocol
Participants satisfying our selection criteria were interviewed remotely for 60-70 minutes, with one researcher facilitating and recording. Each interview had two parts: a general session followed by a task-based session.
General Session. The general session involved 21 open-ended questions to establish the broader context about their decision-making and lasted 45 minutes. We began by exploring the types of decisions participants make, the data they use, and the teams they rely on. Then, we transitioned to discussing their general advanced analytics techniques they use, such as WIA or predictive and prescriptive model-based analytics. Overall questions were grouped into four categories: (1) role and team, (2) business goals, (3) data, tools, and WIA techniques, and (4) ideal tools, as detailed in Table 2.
# Interview Study Questionnaire Role and Team 1 Briefly describe your role in your company? How many people are part of your team and what are their roles? 2 Are there data analysts in your team? Business Goals 3 How often does your daily work require you to answer questions or make decisions to achieve your business objectives? 4 Describe recent business questions that you/your team often try to answer. 5 Approximately how many hours or days do you spend answering each of these questions? 6 Do you answer these questions in one analysis sitting or repeatedly come back to find answers that would eventually help make a decision for your team or company? Data, Tools, and WIA Techniques 7 How often do you/your team use the data to make decisions? 8 Describe the data you or your team utilize in your work. 9 What are the tools or systems (could be commercial, your company products, or self-developed) you generally use to answer such questions? 10 Where does this data come from? 11 Talk us through the analytic process of coming up with strategies or making decisions elaborating on analytic techniques you use in response to one or two specific business questions you shared? 12 Do you consider your current process of using analytics you use efficient? Explain. 13 Do the set of tools you use for performing the analytics you use seem efficient? 14 What are the challenges you face to achieve your objectives? 15 What are the challenges of the set of tools you use? 16 Do you use any advanced analytics like making predictions using the data for your decision-making? Why/Why not? 17 Do you base your predictions on your intuition? Explain. 18 Do you base your predictions on your experience? Explain. 19 Have you ever heard of WIA, predictive analytics, and prescriptive analytics? Can you describe some examples? 20 Do you use any of these analytics that use models or similar techniques (e.g., machine learning, deep learning, statistical modeling, techniques, etc.)? If yes, which ones, why, which tools, expertise level, number of years of experience, examples of how. If no, why not and which techniques would you like to use? If sometimes, why? Ideal Tool 21 If you were to invent the perfect tool to work with data and make decisions with, what would that tool look like? What would it do?
Table 2. Questions that we asked participants during the general session of the interview study. These questions aimed to understand business users’ roles and teams, business goals, data, tools, WIA practices, and desired tools for making data-driven decisions.
Task-Based Session. After the general session, participants spent 15-25 minutes on a task-based session discussing WIA techniques, tools, and challenges in a specific business use case of Marketing Mix Modeling (MMM). Discussion on this specific use case in addition to the general discussion allowed for consistent comparisons across participants.
Marketing Mix Modeling Use Case. Participants, asked to imagine themselves as marketing managers, were tasked with making decisions to achieve the business goal of maximizing TV sales for the next quarter by assessing a historical dataset. The dataset consisted of weekly data on investments made on five advertising channels (SMS, TV, Internet, Radio, Newspaper), economic factors (Demand, Supply, Unit Price, Consumer Confidence Index, Producer Price Index, Consumer Price Index, Gross Rating Points or reach of advertising channels), and company’s sales.
Participants were shown this dataset, sourced from Kaggle (Kaggle, 2022) in CSV format during task briefing. And prompted to discuss advanced analysis techniques and tools they would employ to analyze the data and make decisions to maximize sales. This use case was chosen since it was relatable for many, especially marketing managers, and broad enough for non-marketing participants to apply their own experiences to. Assumptions and validity of both this use case and dataset were verified by our company’s marketing experts, who served as our collaborators. Participants could take notes, if desired.
5. Findings of the Interview Study
We present the interview study findings in three parts: participants’ decision-making context (goals, decisions, and data used), the WIA techniques and tools employed along with specific challenges encountered, and the broader challenges associated with advanced analytics.
5.1. Business Users’ Goals and Decisions
We provide context on participants’ business goals, decision questions, and data used.
Business User KPI Goals Decision Questions Data Marketing click rate, inquiry rate, conversion rate, retention rate which [existing running] campaigns are doing well and why? (P1, P13), how to optimize media mix? (P19), how to move opportunities higher in sales [pipeline]? (P1, P8) search engine optimization (SEO) data (e.g, traffic, impressions, page rank), click rate, customer usage and engagement data, budget data, sales pipeline data, e-mail activity Sales conversion rate, retention rate, churn rate, return of investment what actions must be taken to increase customer retention rate? (P22), how can we leverage previous or present customer to increase sales now? (P9, P14, P18), how do ecological conditions affect churn behavior in comparison to internal factors? (P5) sales pipeline data, revenue and sales data, quote data, product usage data, customer success scores and metrics, campaign data, and external data like shipping costs and delays, geographical data, competitor data Product CSAT score, retention rate, conversion rate, click rate, content management which features of the product (pricing, audience, etc.) and product changes/enhancements are needed in competing markets or for a given geography? (P4, P12), which customer verticals [or segment of a marketplace] are effective? how to increase their market share? how to tap the others? (P2, P4), should we invest in customer needs given the risk? (P2, P11) sales data, market share data, customer success data, CRM data, customer demographics data, customer interaction and engagement data, customer activities, feature usage Operations CSAT score, throughput, productivity, inventory stock how to handle staffing [hiring/management] to reduce costs, maintain productivity, and handle ecological condition changes (e.g., holiday season, inflation, etc.)? (P6, P15), how to cut down wait-times for both products and customers? (P16), how to optimize inventory efficiently? (P16, P17) inventory data (number of products, types of products, count of products, etc.), budget and pricing data, CSAT data, historical data of buy, sell, profits, and sales made from various products, manufacturing or development time, processing time, billing data, time aspects, failure or success data
Table 3. Goals, decision questions, and data sources considered by study participants when making decisions at work.
5.1.1. Goals
Participants, regardless of their role or company, shared a common goal: improving business outcomes and increasing business efficiency to gain a competitive edge over competitors. These outcomes were measured through Key Performance Indicators (KPIs), such as sales, deal closing rate, and retention rate. Decisions focused on maximizing two predominant KPIs—the revenue and sales—while optimizing time and resource constraints. Given the complexity of pursuing these goals, they were broken into smaller, department-specific KPIs. For example, marketing managers P1 and P8 needed to make decisions for the goal of turning prospective customers into real ones, which is measured by the KPI conversion rate. In another example, product managers P4 and P12 were interested in preventing existing customers from leaving, which is measured by the KPI retention rate.
5.1.2. Decisions
To achieve their goals, business users made decisions about actions related to drivers, which were data variables about customers and products. Examples included search engine optimization (SEO) data (e.g., traffic, impressions), product usage data (e.g., feature usage, time spent), and customer activities (e.g., sign-ups, clicks). For instance, to boost conversion rates, decisions included “providing 3 months free trial on the product on signing up for a demo” (which involved the “demo sign up” driver), or “sending out targeted marketing emails showing how the product works for prospective customers’ use cases” (which involved the “marketing emails sent” driver). Similarly, to increase retention rate, plausible decision was “changing placements of product features and highlighting certain features using pop-ups that match customers in a certain geography” (which involved the “feature usage frequency” driver), or “providing tutorial use cases of product mapped to customers’ business niche” (which involved the “customer vertical” driver). More examples are tabulated in Table 3. For a given goal, participants needed to decide both which drivers to act on and how to modify them.
5.2. WIA Techniques Employed by Business Users
Participants used various advanced analytics techniques to understand how independent driver variables relate to dependent KPI business goals and to form hypotheses about these driver variables as decision actions.
5.2.1. Tasks
They followed the underneath set of tasks to understand this relationship:
Frame Questions. With KPI goals in mind (e.g., maximizing sales, optimizing media mix), participants framed decision questions like “what is the relation between the drivers and the KPI?” (P22), “what penetrates” (P15) or “is driving” (P17, P20) the sales, and what to do to “diversify spending [across channels] to minimize diminishing ROI” (P3). Many of these questions aimed at exploring potential scenarios through WIA.
Formulate Hypotheses. Participants often had preconceived hypotheses about how drivers impacted KPI goals, based on their experience, expertise, and external factors. For instance, several participants (P1, P3, P4) shared P8’s expectation that “…expect[ed] newspaper ads to be less effective than internet”.
Translate Hypotheses into Models. Participants refined their hypotheses into simple quantitative models. For example, a marketer (P19) aimed to optimize impression rate (i.e., number of customers reached or tapped by a marketing action) by understanding how marketing channels (e.g., ads, website) contributed to the impression rate KPI by weighing how much each channel was contributing towards one impression. They came up with a rough hypothesis that a potential customer looked at two paid media ads and visited the website, it would result in a complete impression, which was translated into a simple linear model (e.g., 2 $\times$ paid media ads + 1 $\times$ website visit = 1 impression.
Revise Models with Data. Participants frequently adjusted their models by varying parameters based on intuition and trial-and-error. Data played an essential part, and participants relied on existing business intelligence (BI) reports, visualizations, and raw data to refine their models. For example, after reviewing more data and playing with more models, P19 updated their formula to include video completion rate to lead to a more successful impression, changing the model formula to 2 $\times$ ads + 1 $\times$ website visit + 1 $\times$ video completion rate = 1 impression instead. This highlights the need for business users to test various driver-action scenarios (Dimara et al., 2021; Wang et al., 2019) and observe it’s impacts on a given KPI before making strategic decisions.
Explore Multiple Scenarios through WIA Techniques. Overall, all participants explored multiple what-if scenarios. Nearly all participants (20 of 22) manipulated driver or KPI values to observe their effects on other variables in the model, while the remaining participants expressed interest in using such advanced analytics. Examples of the varied WIA techniques discussed in our interviews are shown in Figure 3 and described in order of popularity:
Figure 3. What-if analysis techniques shared by the participants for informing their decisions using data.
  • Sensitivity Analysis: Maximum participants, 20 out of 22, wanted to observe how changes in driver values (e.g., advertising spend across different channels) affect key performance indicators (KPIs) like sales. For example, P7 wanted to know if “they should do something to the unit price so the return [sales] is better?” Such analysis is commonly referred to as sensitivity analysis, where users explore how altering input variables influences output variables (Rappaport, 1967).
    When using spreadsheets, participants created models or formulas for KPIs, as described in earlier tasks and manually adjusted the driver values to observe how recalculated KPI formulas to avail this technique. They also tweaked formulas to simulate and experiment with different scenarios to explore the impact of changes on various driver combinations, helping guide decisions on next steps. While a few participants (2 out of 22) mentioned existing tools like Google Analytics, Power BI, and Salesforce that offer text-based insights or narratives of few scenarios, they felt these tools were too static. They preferred more dynamic, interactive tools that allow for real-time manipulation of drivers, enabling them to test multiple scenarios and see the immediate impact on KPIs.
  • Driver Importance Analysis: 17 out of 22 participants wanted to determine which drivers (e.g., advertising channels) were most significantly influencing the sales KPI, often known as key influencers or driver importance analysis. Currently, participants relied on manually formulating hypotheses about how drivers related to sales KPIs and performing sensitivity analysis by adjusting driver values from far negative to far positive values to observe changes in sales.
    Participants typically used trial-and-error, adjusting driver values (far negative to far positive) to see which changes had the greatest impact on sales. This manual approach relied heavily on domain knowledge and self-constructed formulas for KPIs, making it difficult to fully understand driver-KPI relationships. Participants acknowledged this limitation, as their approach relied heavily on domain knowledge and guesswork. Although tools like Salesforce Einstein offer a feature that ranks drivers by importance, none of the participants mentioned using it. One participant (P8) noted seeing a similar feature in Power BI’s waterfall chart, which is also seen in Einstein, but found it hard to interpret when quick, actionable insights were needed.
  • Segmentation Analysis: Half the participants expressed interest in segmentation analysis, that involved understanding the needs and behaviors of various drilled-down groups of data such as different “regions” or “user types”, as shared by P10, P14, P18. Participants used this technique to tailor their strategies for distinct data segments. To do this, participants manually segmented data in spreadsheets, copying and pasting slices and dices into separate sheets for individual analysis and comparison.
    No existing tools were mentioned as helpful in this context. While platforms like Tableau and Salesforce Einstein provide text-based insights on certain segments, participants noted them to be limited in scope and often one-off that may not always be actionable or relevant to their specific needs. Additionally, if business requirements for a segment change, these could not be reflected in the underneath data to simulate its impact effectively.
  • Correlation Analysis and Outlier Analysis: 10 of 22 participants expressed the need for learning correlations between different drivers and their KPI to help identify whether, in which direction, and to what extent changes in the drivers are associated with changes in the KPI. To make it easier to digest it, they also wanted to learn a summary of the impacts.
    About three to four participants discussed conducting correlation analysis between drivers to learn the trend of a range of changes made to the drivers on the KPI. They mentioned creating heatmaps to visualize correlations, but also wondered how to justify shifting results to other team members when models were changed.
    Additionally, about three to four participants also mentioned support in detecting anomalies in the data using outlier analysis, but only mentioned currently relying on manual inspections for detecting outliers.
  • Comparison Analysis: Half the number of participants were also eager in comparison analysis where they could actively compare two or more of any of the advanced analyses, but found this to be difficult and confusing currently.
  • Goal Seeking Analysis: Few participants came up with a set of driver actions to trigger the desired changes in their KPI goals. P7, for example shared “…send to different email, …change subject line, …colors, … time” actions to run an email promotional campaign to get “1000 more downloads”. This aligns with the basic concept of goal-seeking analysis.
    One participant (P11) referenced Excel’s GOAL SEEK macro, which predicts the necessary driver values to meet a target KPI but shed light on high expertise levels to effectively use it, while another participant, P21, mentioned not using “it enough or have the time enough to use it… but can already see the value…”.
  • Constrained Analysis: Participants also explicitly shared budget and resource constraints while making decisions. For instance, a participant P2 shared they “obviously have a limited budget about what campaigns to do…” and wanted to know if there were means to include these constraints during analysis, rather than relying on manual data adjustments. This is an example of constrained analysis.
Formulate Initial Actions. After exploring multiple scenarios, participants devised actionable decisions based on their analysis. For example, P19’s revised model led to doubling resources on paid media ads and instructing sales managers to push prospective customers to visit the website and see the complete video explicitly during interactions.
5.2.2. Tools
All participants shared mainly using spreadsheet applications, primarily Excel for the WIA techniques. Non-technical business users relied on manual methods, explicitly creating formulas and running calculations or using operations like filter and pivot to manually slicing and dicing the raw data in different ways, while technical users employed programming tools and machine learning models (Kandel et al., 2012; Alspaugh et al., 2019; Crisan and Fiore-Gartland, 2021). Only one participant, P10, reported using advanced functions (e.g., LOOKUP, VLOOKUP to retrieve data from large tables) and features (e.g., GOAL SEEK) but highlighted a lack of time to set up and use.
Participants also shared creating charts, but only to observe patterns (e.g., seasonal trends), behaviors (e.g., sales best in the west region), or anomalies (e.g., sudden sales drop every 5 years). Only 5 of 22 visualized their data in dashboards (static and interactive) using a variety of BI tools like Power BI, Salesforce, and Tableau, while 17 others relied on creating Excel charts. Those with access to technical data teams (8 out of 22) or other third-party application tools (e.g., Google Analytics, Marketo, and Pendo) used team or application-generated reports that were generated generally on a daily, weekly, or monthly time-frame. However, those who could not get customized reports easily often struggled with using the BI tools on their own.
To make data-driven decisions, most participants (17 out of 22) mentioned performing complex analyses described earlier (Section 5.2.1), however executing them heavily using manual and simple processes like updating tables and utilizing formulas. The descriptions provided by other participants implied a similar need for advanced analytics and collectively all participants called for more sophisticated means to conduct them easily.
5.2.3. Challenges
We identified several broader challenges faced by business users in leveraging advanced analytics for making data-driven decisions, detailed as follows:
Current Advanced Analytics Use Rudimentary Methods. Challenges with advanced analytics are amplified for business users due to their lack of formal technical training. As discussed, they often rely heavily on spreadsheet tools for performing advanced analytics, which, while “super powerful” (P11), come with drawbacks of “need[ing] a lot of experience with [more complex] macros, etc. which not everyone has[d]” (P11), making even simple analyses inefficient, time-consuming, and prone to errors. For example, P11 recounted frustration dealing with extremely large boilerplate Excel sheets made 20 years ago by some individual and how “absolutely non-replicable the process is and how the person who developed it is still contacted at times to understand some niches”.
Participants also appreciated descriptive and exploratory features in BI tools. For example, visual data exploration capabilities of Tableau were praised, calling them “spectacular” (P5) and “Excel on steroids!” (P4). However, these tools were mainly used for exploring patterns in data lacked capabilities to generate and confirm participant’s hypotheses, necessary to inform data-driven decisions. Further, participants also criticized these tools for being static or minimally interactive, offering only basic data statistics, percentage of goals achieved, and limited drill-down scenarios.
Even for those having advanced analyses features, participants disliked them being cumbersome, noting they “need[ed] to have a developer support to use it… and extensive training” (P4) and describing them as “clunky… [with the] filters, …not very user friendly compared to others” (P5). An employee at one of the BI tool companies also admitted that they “are not proud of not using [their own tool], but it does have a steep learning curve” (P21). Participants also highlighted the lack of state-saving features, limiting their usefulness for iterative analyses. Participants needed more dynamic and interactive tools to simulate various scenarios, test hypotheses, and understand driver-KPI relationships effectively, that allowed for more control and proactive support for advanced analytics techniques.
Business Users Conduct Advanced Analytics without Relying on Technical Users. All participants reported a shortage of analysts in companies, regardless of their size, to handle business users’ requests. They mentioned that data teams also have limited bandwidth and prioritize company-wide initiatives (e.g., “developing models for churn rate analysis” (P4, P5), “customer ordering patterns” (P6), “staff hiring models” (P5, P6)) and automation over their needs of analyses queries. Even when analysts were available, participants expressed frustration due to analysts’ lack of domain knowledge and inefficient communication which made the collaboration “a lot of added work [inconvenience] in itself” (P3). Additionally, under the pressure to make quick decisions, business users often did not have the time or patience to engage with data analysts or scientists, spending mostly 1-4 hours (20 out of 22) and at times 4-6 hours (2 out of 22) daily on analysis themselves without relying on them. These findings highlight a gap between ideal collaborative practices and the real-world challenges of achieving effective communication and understanding and goal alignment between business users and data analysts (Davenport et al., 2010; Provost and Fawcett, 2013).
Unique Challenges Exist Beyond Data Analysis. Before delving into advanced analytics to understand driver-KPI relationships, business users often undertake tasks similar to data analysts (Kandel et al., 2012; Alspaugh et al., 2019; Kandogan et al., 2014), such as acquiring, cleaning, and enriching data and face similar challenges of integrating data from various sources, ensuring data quality, etc. but more aggravated since they are not technical users. When they conduct advanced analytics independently, without the support of technical data scientists or analysts, they face additional challenges. For example, they must balance various business factors, such as budget constraints, economic principles (e.g., the impact of price increases on sales), market conditions (e.g., COVID, interest rates), and social considerations (e.g., employee work-life balance), all while under pressure to make accurate and actionable decisions. Also, participants perform analyses continuously, where they experiment with hypothetical scenarios, refine hypotheses, formulate decision actions, share them with stakeholders, and conduct A/B testing over smaller groups of customers over multiple iterations to execute decision actions that eventually help achieve their goals. Consequently, business users’ analyses to make data-driven decisions are inherently more complex, extending beyond traditional exploratory data analysis.
6. Study II: Follow-Up Task-Based Study
The first interview study informed various WIA techniques business users employed and the current difficulties they faced with them. In the task-based study, we implemented four representative WIA techniques identified in the previous study in a visual analytics prototype as a probe to gather participants’ hands-on experience and feedback on the benefits and enhancement opportunities of WIA techniques for making data-driven decisions.
6.1. Method
We outline the methodology of our task-based study, followed by a description of the WIA techniques implemented in a visual analytics prototype to use as a probe.
6.1.1. Protocol
We conducted a 60-minute remote session with the same participants for the task-based study, including: (1) a tutorial, (2) a re-introduction to the marketing mix modeling (MMM) case and task, (3) a hands-on session with the WIA techniques, (4) open feedback, and (5) a post-study questionnaire, as shown in Figure 2.
Tutorial Viewing. We provided a 15-minute pre-recorded video of the visual analytics prototype to acquaint participants with four representative WIA techniques implemented therein: (1) driver importance analysis, (2) sensitivity analysis, (3) goal-seeking analysis, and (4) constrained analysis, as shown in Figure 4. The demo used another common business case, deal closing analysis, which sales managers use to increase customer acquisition rate by analyzing prospective customer’s interaction data with the product and the organization.
Re-Introduction to the MMM Business Use Case and Task. Participants were refreshed with the MMM use case and dataset to set context for the hands-on session. They were briefed on the same task of maximizing TV sales, but informed to use the representative WIA techniques. This lasted 5 minutes and participants could take notes if desired.
# Task-Based Study Sub-Tasks and Open-Feedback Questionnaire Sub-Tasks 1 Using driver importance analysis technique, what are the top three and least three drivers that influence the sales KPI? 2 Using sensitivity analysis technique, which advertising channels will you invest in in order of most to least to increase the sales KPI? 3 Using the sensitivity analysis technique, what is the sales achieved if you increase the Demand driver by 6% and the Supply driver by 3%? What can you say about the demand-supply relationship? 4 Using the sensitivity analysis technique, find the percentage increase in the Unit Price driver needed to achieve the same or more sales if the Demand and Supply drivers don’t change. 5 Using the summary analysis technique, discuss the trend observed of perturbing the non-advertising channel expenses drivers between a range of -100% to 700% and step size of 100% on the sales? Do you expect this behavior? 6 Using the summary analysis technique, discuss the trend observed of perturbing the advertising channel expenses. Which advertising channels would you invest or not invest in to increase the sales of the TV? Do you expect this behavior? 7 Using the goal seeking technique, what is the maximum TV sales you can achieve with the original boundary conditions? 8 What is the sales uplift that you can achieve if you only optimize the advertising channel drivers and the Unit Price driver? 9 How much will the sales increase if you have constraints of \$2-\$115 on the Unit Price driver? 10 Can you achieve a target sales of \$980 million? If yes, what is the driver that needs the maximum uplift? Do you think this is feasible? Open-Feedback Questionnaire 1 Were these specific tasks relatable to the ones you would try to answer to achieve your goal of increasing sales or any other decisions that you take in your own work? 2 Do you understand the key driving advertising channels behind sales? 3 Do you understand the relationship between the advertising channel expenses and sales? 4 How open would you be to using such WIA techniques for answering your own business questions? 5 What would be some other advanced analytics techniques that would be helpful for you to achieve your goal? 6 Were you able to reach a decision on investments needed to achieve maximum sales? Explain. 7 How confident are you in the decision you reached? Explain your reasoning. 8 What are your concerns about your inferred decision? Explain your reasoning. 9 What were the challenges you faced to achieve the goal of increasing the sales of the TV company? 10 Do you consider the analysis process you currently used of using the WIA efficient? Explain your reasoning.
Table 4. In the hands-on session of the task-based study, we asked participants to complete a set of sub-tasks (bottom) requiring them to use what-if analysis techniques implemented in an interactive visual analytics prototype. In the subsequent open feedback session, we asked participants to answer a set of questions (bottom) in a questionnaire, eliciting their qualitative feedback based on their task completion experience.
Hands-On Session. Participants accessed the prototype via an AWS-hosted application link. Based on a pilot study with 3 participants (not included in the 22), we gave each participant 25 minutes and broke down the task into 10 manageable sub-tasks, as tabulated in Table 4). They were instructed to think aloud, given a document detailing each sub-task, and received assistance as needed. Participants were additionally prompted to share how each technique could be applied to their own decision-making use cases.
Open Feedback. Following the hands-on session, participants provided 10 minutes of feedback on the WIA techniques’ utility, benefits, concerns, and any additional analytics they wanted. This session also gathered insights on their trust, confidence, and perceived effectiveness of the techniques. Any missed feedback was collected using the open-feedback questionnaire (Table 4).
Post-Study Questionnaire. In the final 5 minutes, participants completed a post-study questionnaire (Anonymous, 2024c) comparing their own analyses practices with those supported by WIA. Details are in the supplementary materials (Anonymous, 2024a), and all sessions were recorded for further analysis.
6.2. Representative WIA Techniques Implemented into a Visual Analytics Prototype to Use as a Probe
We implemented four representative WIA techniques identified in the previous study within a visual analytics prototype to use as a probe in this task-based study. Although we focused on a subset to avoid overwhelming users, the chosen techniques covered a well-rounded set of viable techniques to help business users make data-driven decisions.
To ensure a comprehensive experience and enable making actionable decisions, we implemented techniques from across the pool mentioned by participants in the interview study including less popular ones. However, we made it clear that these techniques were only representative and the visual analytics prototype was only a probe to capture users’ practical and critical feedback, not an exhaustive set or tool. Throughout the study, participants were also encouraged to consider other advanced analytics techniques that could further enhance their decision-making.
We demonstrate the techniques are using the MMM business cases but they are applicable to other use cases described by business users like increasing customer retention, decreasing churn rate, and optimizing inventory. For two exemplar demonstrations, please refer to the supplementary materials (Anonymous, 2024a). We describe the representative WIA techniques in an order reflecting how business users typically approach understanding driver-KPI relationships. Nonetheless, findings from our study revealed that participants often preferred to use them in any order and iteratively.
Figure 4. Interface showing the four representative WIA techniques implemented in a visual analytics prototype, used as a probe to capture participants’ practical use of WIA techniques for making decisions in the MMM business use case during the task-based study: (1) driver importance analysis bar chart (A) with toggle-able drivers (B), (2) sensitivity analysis perturbation pane (C), bar chart (D), and uplift pane (E), along with its variant of summary analysis (G) with bounding pant (F), (3) goal-seeking analysis prediction (H), (4) constrained analysis options and constraints pane (I), and prediction (J), and the $R2$ accuracy score pane (K).
Driver Importance Analysis. This technique helped understand the relative importance of drivers in predicting the sales KPI, such as the top three advertising channels to prioritize. We implemented this technique as an ordered list of drivers (Figure 2B), accompanied by a bar chart (Figure 2A) showing the computed importance values, offering a simple to read, systematic, and insightful approach compared to approaches seen in existing tools. Using this, users quickly learnt that the top three and bottom three drivers for the sales KPI were CCI, PPI, and Supply, and Newspaper Ad Expenses, GRP of Radio, and GRP of Newspaper respectively.
Sensitivity Analysis. This technique helped observe the effects of adjusting driver values, such as increasing or decreasing expenses on different channels, on the sales KPI. This was provided in the form of coordinated views, comprising a perturbations pane (Figure 4C), a bar chart (Figure 4D), and an uplift pane (Figure 4E). The average sales observed on the original data was 96 million dollars (blue bar); upon perturbing the “supply” driver by 3% and “demand” driver by 5%, the predicted sales dynamically shows an increase by 3.63% (equivalent to 3.5 million dollars) (green texts), making the overall predicted sales 100 million dollars (yellow bar).
Summary analysis (Figure 4G) allowed users to view multiple sensitivity scenarios at once, that is aided them in comprehending the sales KPI trend for when individual driver values underwent perturbation across a specified range of bounds (Figure 4F). A small multiples chart was used, with each multiple displaying the trend of the sales KPI impact when each advertising channel expense was varied from -100% (lower bound) to 500% (upper bound), with a step size of 100%, revealing significant sales increases for SMS and Radio channels, a slight rise for the Internet channel, and declines for TV and Newspaper channels.
Goal-Seeking Analysis. The technique helped users learn predictions of the required driver values to achieve an optimum or user-desired aggregate KPI value, like to directly identify the actions needed to maximize the sales KPI. We implement it iteractively (Figure 4H), where users chose their optimization goal (maximize, minimize, or set a target goal), observed the pre-specified constraints on all drivers to initially be optimized between 0% to 300%, and ran the analysis. On maximizing the sales KPI goal, a prediction of 549.07 million dollars sales could be achieved (470.11% uplift) along with the driver values that enabled this prediction.
Constrained Analysis. This technique incorporated user-defined business constraints (e.g., boundary, equality or inequality conditions) on one or more drivers and run goal-seeking analysis to get driver values satisfying these constraints. For instance, when using the previous goal-seeking analysis technique, participants observed that in order to maximize sales, the “unit price” driver needed to increase by 297.81% or 74000 dollars, which customers would not practically pay. In response, they could restrict the “unit price” driver to a 10% to 15% change in the constraints pane (Figure 4I) which reduced the sales prediction from 549.07 million to 277.84 million dollars, but with considering “unit price” driver prediction to be 14.08% within constraints, and yet showing a significant uplift of 188.49% while predicting the “unit price” to be 14.08% (Figure 4J) within given limits.
Each technique also included accuracy ($R^{2}$ measure) or error scores (error rate) of the underlying prediction models, with the MMM use case showing a fair accuracy of 0.82 (Figure 4K).
7. Findings of the Follow-Up Task-Based Study
We present our findings by first reporting on participant’s experiences of using the representative WIA techniques for the MMM business use case. Then, we discuss the potential benefits and opportunities for enhancement they identified regarding the use of advanced analytics for making data-driven decisions.
7.1. Insights From Using Representative WIA Techniques for the MMM Business Use Case
Figure 5. Ratings of the what-if analysis techniques from most to least useful. Participants found sensitivity analysis to be the most useful technique.
Feedback from both the open-feedback session and post-study questionnaire showed unanimous agreement that WIA techniques implemented in the visual analytics prototype as a probe were “helpful” (P2, P10, P12, P17, P20), “interesting” (P3, P8), “powerful” (P13, P16), and quick to “set [them] in the right direction” (P7, P17) for decision-making to maximize the sales KPI. All participants comfortably completed all 10 sub-tasks within the allotted time, finding them simple and easy to use, and made similar decisions, such as increasing the unit price of TVs, boosting SMS advertising spending, and cutting newspaper ad spending. They also mentioned that using WIA techniques was “better than sitting looking at the data” (P9) that they did in their existing workflows. All but one participant were excited to apply the techniques to their own business use cases, to “play around with own data [since it would be] super helpful when setting goals for the quarter” (P3) and learn data behaviors like “how it can relate to their business” (P6), “how it would be useful for drilled-down segments [subsets of their data]” (P4). The remainder participant, an operations manager, found the techniques to be “okay and wanted [them] to be more dummy-proof” (P15) suggesting their utility, but also the need for more greater customization towards their domain.
As for the usefulness of the techniques, participants collectively ranked them all as useful (Figure 5), with sensitivity analysis rated highest (most saturated stack), followed by constrained analysis, goal-seeking analysis, and driver importance analysis (least saturated stack). The goal-seeking technique required the most guidance (6 out of 22 participants), despite aligning well with their decision-making needs, highlighting the importance for explicit guidance, demonstrations, and examples for their practical use. While participants also initially struggled with interpreting the small multiples chart expressing summary analysis, they found it valuable for reducing cognitive load in experimenting with multiple what-if scenarios. However, they stressed the need for simpler visuals when presenting such advanced analyses to executive stakeholders.
Further, participants expressed a keen interest in additional advanced analytics techniques that they also mentioned in the first interview study. They sought automated features with robust drill-down capabilities, enabling segmentation analysis to help “how different geographical locations are doing?” (P14), “realize how the data behaves for different verticals of customers?” (P18), or “how to leverage existing customers and products doing well for those not performing?” (P21) They also wanted techniques to uncover correlations between drivers and KPIs and facilitate easier scenario comparisons. Participants requested enhancements to the WIA techniques, like seamlessly integrating results from one technique into another (e.g., plugging in the predictions from goal-seeking analysis back into sensitivity analysis (P21)) for continuing the analysis, and the ability to save analysis results for future comparisons (P1, P8, P14).
7.2. Benefits of Using WIA in Making Data-Driven Decisions
We report three key benefits shared by participants in using WIA techniques for making their data-driven decisions.
7.2.1. Helps Accelerate Decision-Making
Participants appreciated the ability of WIA to provide a “holistic view” (P13) of driver-KPI relationships, unlike the currently used manual approaches of attaining it using spreadsheet-like tools. More importantly, participants were happy that they could quickly observe the predictions for many scenarios encompassing various combination of drivers all together without the need for external expertise from data analysts or teams. Consequently, all participants shared that WIA would speed up their decision-making and “increase [their] efficiency” (P6). Participants were also pleasantly surprised by the capabilities of the representative techniques since they may have seen “something similar but not anywhere close to these [representative techniques]… to customize them or make your own if you want anything like these [representative techniques]…” (P11). As such, participants emphasized the transformative potential of incorporating advanced analytics in streamlining and expediting their decision-making.
7.2.2. Increases Confidence in Decisions
Participants articulated that WIA helped them move away from from guesswork and trial-and-error approaches, thereby bolstering their confidence in their decision-making (Elhamdadi et al., 2022), as it narrows down the vast possibilities of hypotheses to those observed in the data. One participant P1, for instance, happily shared how WIA can “help understand where to put more effort!” Moreover, participants emphasized that leveraging advanced analytics techniques, particularly interactively as demonstrated in our study, could help them think and test out a lot more corner cases and scenarios that they might miss in traditional manual analysis. Further, participants shared how they could play with many more drivers to observe their consequences on the business KPI and be able to understand the market, product, and customer behavior more widely and better. All these reasons combined led the participants to excitedly share how they can build a strong narrative or story supporting their decision actions. A participant P10, for example, expressed how WIA will enable “everyone in the company [to] be on the same page about a decision… no one’s decision will be overwritten”. Therefore, participants showed promise in WIA enabling them to be confident in their decisions and ready when communicating with the stakeholders.
7.2.3. Streamlines Decision-Making and Enhances Teamwork
Almost all participants agreed that their existing analysis processes could not be replicated easily to justify their decision actions. For instance, multiple participants commented that their current decision actions were based on “trial and error” (P3, P14, P19, P21), which they often deemed to be inefficient and unpredictable, making it challenging to rely on. Recognizing the limitations of their current practices, participants saw the potential of WIA to streamline decision processes, making them more accessible, reproducible, and actionable. This consequently allowed for “consistent” (P3, P10) (Kandel et al., 2012) and “smoother” (P22) decision-making. Further, in addition to WIA techniques helping them in their own analysis, participants also added that it will help facilitate collaborative analyses for decision-making envisioning a scenario where stakeholders could engage in decision reasoning collectively. For example, one participant P4 mentioned how they could “take [this] to their job today without the support of IT… also give it to other team members”. Therefore, participants explained that advanced analytics techniques would help them make robust decisions as business users could get external perspectives and benefits from others’ experiences and expertise through standardized and collaborative analysis processes.
7.3. Opportunities in Enhancing the Use of WIA for Making Data-Driven Decisions
After learning the benefits of WIA, we outline three opportunities raised by our participants to enhance the utilization of WIA techniques for making data-driven decisions.
7.3.1. Currently Inaccessible to Business Users
Many participants like P1, P7, P8 mentioned that even though they had heard of features similar to some representative WIA techniques in existing commercial BI tools (e.g., key influencers in Power BI, importance and correlation analysis in Einstein, etc.), they had “no idea or not know too much [about them] or barely used them”. Upon further discussion, participants revealed that high costs associated with such tools rendered them inaccessible. Additionally participants voiced concerns about the complexity of these tools, which typically offered a plethora of models and parameters to navigate, or deemed them to be too general and inefficient to use for their own use cases. The overwhelming number of features too made it challenging for users to effectively utilize the tools, exacerbated by automatically generated results with a lack of clear explanations for them to draw conclusions from. Moreover, participants cautioned the requirement of technical expertise to utilize these tools effectively, a skill set that many business users did not possess. Furthermore, business users also lacked the time and incentives to learn or to keep up with such technologies, which are rapidly evolving.
Despite their enthusiasm for adopting WIA in their work, it seems that tools are not currently designed to meet their needs. The vast space of possible scenarios seems to be a major problem in particular. P15, for instance, frustratingly mentioned how they have “…to [currently] commit to just solving the first thing… that is going to be the most bang for the buck…. [because]… have 900 problems happening all at once…”. Such expressions inform us that business users have no time to experiment with all possible scenarios but need flexibility to quickly identify the first set of actionable decisions that yield favorable business outcomes and play with them to account for constraints, a concept akin to “satisficing” (Simon, 1960), suggesting that more complex analyses atop WIA, like recommendations or ranking of WIA scenarios, could help.
7.3.2. Difficulties in Interpreting and Understanding Predictions
While participants expressed enthusiasm for utilizing WIA techniques for decision-making, they were clear that such analytics alone cannot make the decisions. Their domain expertise remains critical in strategizing decisions (Wang et al., 2019). For instance, P4 pointed to how “common human understanding may not translate to the model without [their] external involvement”, giving an example where an increase in unit price beyond a threshold will decrease demand, thereby decreasing sales rather than increasing it, a nuance that may not be fully captured by models. Participants emphasized the need for transparency in how predictions were generated and wanted clarification on when and to what extent to adopt predictions. For instance, they wanted to know how to handle “unrealistic [predictions]” (P5, P6, P22), or when “so many factors [or decision paths]” (P2) are shown to them, and how to “[include] factors that cannot be captured in the data” (P2, P18).
Another recurring challenge was interpreting the accuracy of predictions (low v/s fair v/s high) and translating this into confidence in next steps (Elhamdadi et al., 2022). For example, should low-scored models be treated as worst-case scenarios and perform additional manual analysis to make decisions? Or should they take medium scored models as baseline and expect risks in lower scored models? While senior, more experienced business users understood that predictions should be taken with caution, they emphasized the importance of training junior users not to follow predictions blindly. Addressing these concerns requires WIA tools to foster deeper transparency and provide interpretable predictions that can help enhance business users’ trust in the predictions.
7.3.3. Lack of Support for Data Preparation and External Constraints
Participants in our study shared various challenges with data gathering, consolidation, featurization (e.g., curating new hypothesized drivers for KPI goals), aggregation from multiple sources, and cleaning it to work around missing, inaccurate, and outlier data. These operations are already hard for expert analysts (Dimara et al., 2021; Crisan and Fiore-Gartland, 2021; Kandel et al., 2012; Alspaugh et al., 2019), and our study confirms that it can be all the more daunting for business users. For example, the participants raised some fundamental issues with adding qualitative data (procured over calls, meetings, chats, and reviews) as drivers into models, treating connected drivers (e.g., calls and meetings), and resolving ambiguities about the drivers or KPIs as they could mean different things to different teams at times. Therefore, participants suggested co-designing and tight integration of WIA techniques with data wrangling and preparation systems.
Although participants valued techniques like goal-seeking and constrained analysis to incorporate their domain expertise, they worried about the mechanisms “on setting the right constraints, coming up with them, how to set them right and optimize on that” (P5). On a related note, some participants expressed concerns over not being aware of ideal conditions for the team and company, thereby gauging the constraints either conservatively or over-confidently. For example, P5 mentioned only knowing that they “had five million dollars for advertising… [and] can set only overall constraint over all channels” making it challenging to set constraints on each of the channels. Therefore, participants requested features to help set constraints on the data more effectively.
Further, participants also wanted to factor in economical (e.g., inflation, fed rate hike) and social (e.g., COVID, war, street protests) constraints into advanced analytics (P6, P9, P10, P15) since they had very significant impacts on businesses. Participants often dealt with these constraints for making decisions by turning to their intuition (Wang et al., 2019) and domain expertise. As a consequence, participants indicated that WIA techniques need to be enhanced to increase their reliability so that it can actively be used “as an adjunct” (P8) to inform their decisions. Therefore, future tools need to focus on feedback mechanisms between business users and advanced analytics which will allow users’ expertise as well as external constraints to be easily integrated into predictions (Zhang et al., 2020; Honeycutt et al., 2020).
8. Discussion
Our two-part study confirms that business users are eager to independently conduct data-driven advanced analytics to inform their business decisions. Based on our findings, we discuss design recommendations for future business analytics systems to better support these advanced analytics needs.
8.1. Enhance Interpretation and Confidence
Our study reveals that business users are far from naive; they possess strong domain expertise and common sense, enabling them to critically assess predictions from automated tools. All participants indicated that WIA predictions cannot be followed blindly (Wang et al., 2019; Kaur et al., 2024), recognizing that even the most advanced models are based on limited data. This data often fails to capture all relevant constraints and domain-specific knowledge, resulting in predictions that may not be entirely accurate. However, participants emphasized the value of advanced analytics in complementing their decision-making, supporting their actions with data-driven evidence.
To foster trust and ensure users feel confident using data-driven analyses, systems must not only be transparent in communicating confidence levels in predictions but also offer actionable steps to calibrate these confidence levels. Participants suggested that model accuracy scores or error metrics should be corroborated with risk factors that communicate potential negative outcomes, such as potential financial losses or the emergence of new competitors or products. Moreover, users desired guidance on next steps based on risk levels (e.g., a medium risk factor might suggest gathering more data and re-running the analysis). Recent work has explored general users’ perceptions of trust and performance metrics (Langer et al., 2022; Zhang et al., 2020; Honeycutt et al., 2020; Bhattacharya et al., 2024) Conducting similar studies specifically with business users can reveal other factors that enhance their trust in model predictions and, in turn, guide the design of advanced analytics systems that better meet their needs. Additionally, future research should investigate the requirements of technically proficient business users, focusing on analyses features that optimize their decision-making. This opens up a substantial opportunity for interdisciplinary collaboration to refine WIA techniques, making them more relevant and effective for business contexts.
8.2. Foster Communication and Team Collaboration
Participants consistently emphasized the importance of WIA in empowering business users to justify their decisions with data, providing them with a stronger voice when navigating organizational hierarchies. As business decisions often require approval from executive stakeholders, it is crucial for users to effectively communicate the rationale behind the proposed decision action predictions of advanced analytics techniques, especially since neither the business users nor the executives are typically data science or machine learning experts. Since business decisions often require approval from executive stakeholders, it is essential for users to clearly communicate the rationale behind decision-making predictions generated by advanced analytics techniques. Additionally, this must be done with the understanding that both business users and executives usually lack expertise in data science or machine learning. This presents a clear opportunity to rethink traditional tools used. For example, current dashboards typically display static summaries, such as the percentage of goals achieved or fixed drill-down paths into data segments (Tory et al., 2021; Sarikaya et al., 2018). These high-level overviews, while useful, do not provide the flexibility needed for decision-making contexts (Oral et al., 2023). Similarly, spreadsheet tables can be cumbersome and confusing (Bartram et al., 2021), and individual presentations prepared for every meeting require substantial manual effort. By contrast, interactive visual tools accommodating advanced analytics could revolutionize these processes. Instead of static, one-way communication, such tools could allow users to dynamically explore various scenarios and their impacts on KPIs. For example, during a meeting, a business user could adjust variables in real-time to demonstrate how different marketing strategies might affect sales forecasts. These tools could facilitate live discussions and decision-making, similar to how participants could see the immediate outcomes of various strategies. Moreover, these tools could be designed to quickly incorporate feedback from executives or other team members during the meeting itself (Spitzeck and Hansen, 2010), rather than requiring additional time to go back, rerun analyses, and return with new findings. For instance, if an executive recalls a similar scenario from past experience or suggests adjusting a particular variable due to new constraints, the tool could instantly update the analysis to reflect that feedback. This kind of interactive, iterative process would not only save time but also foster collaboration, ensuring that decisions are informed by both data and experiential knowledge without the delays that are common in current workflows. Further, emerging technologies such as natural language interfaces and interactive storytelling platforms (Sun et al., 2024; He et al., 2024) must also be leveraged in such tools to go beyond merely presenting predictive results to also additionally explain, modify, and contextualize the predictions from advanced analytics.
8.3. Scale WIA with Decision Management Support
While WIA techniques allow for rapid computation of what-if scenarios and offer insights into various decision paths, participants observed that the number of scenarios and corresponding decision paths can grow exponentially as the number of driver variables and constraints increase. Hence, managing and reusing these analyses at scale presents a formidable challenge, especially as users increasingly gain access to real-time data. To prevent business users from becoming overwhelmed or, as one participant described, from “getting lost in the sauce” (P9), it is essential to develop robust decision management systems that can effectively monitor, track, and log users’ analyses. These systems must be capable of handling large volumes of data and complex decision paths without compromising usability. Though such provenance-based systems have been widely developed for software developers (Schreiber et al., 2021; Packer et al., 2019), data analysts (Cutler et al., 2020; Madanagopal et al., 2019; Ragan et al., 2015), data scientists (Namaki et al., 2020; Souza et al., 2022; Kumar et al., 2016), and researchers (Jun et al., 2019, 2022a, 2022b; Gu et al., 2023) for various analyses (e.g., visual analysis, statistical analysis, software development, etc.), there is gap in supporting non-technical domain users like business professionals. These users require more accessible and user-friendly solutions tailored specifically for analyses that help inform data-driven decisions.
Further, decision management support systems can significantly enhance incorporation of WIA by enabling tools for ranking, comparing, and recommending optimized decision paths. Our study indicates that such integration, as suggested by business users, will significantly improve their ability to leverage advanced analytics for truly data-driven decision-making. However, to maximize the success of these systems, it is essential to focus on usability and accessibility. For instance, systems supporting multitude of models, tons of features, and complex visualizations and workflows or programming language-based tools will not meet the needs of business users. While natural-language, query-based features powered by LLMs may be very useful, as seen in software testing (Santos et al., 2024) and cyber-physical systems (Elmaaroufi et al., 2024). Therefore, besides the vast scope for machine learning modeling and optimization research, there is a huge opportunity for user experience and data and visual analytics communities to contribute to the development of scalable, user-friendly decision management systems that effectively support advanced analyses for decision-making.
8.4. Limitations
We acknowledge the limitations of our studies. First, we are aware of the limited participant pool, representation of business users across four departments. Despite this, similar number of participants have been studied in previous visualization and HCI research studies (Kandel et al., 2012; Dimara et al., 2021; Alspaugh et al., 2019), and we chose few participants and covered multiple departments to ensure in-depth interviews and not have an extremely narrow collection of empirical data.
Second, our findings in the interview study are based on participants’ talk-through and self-reported experiences rather than actual workflows from their own use cases. Time constraints and the need for understanding their advanced analyses techniques comprehensively led us to choose verbal articulation over demonstration. We addressed this by asking them to talk us through a specific but common marketing use case that was familiar to participants. Additionally, in the follow-up task-based study, participant’s hands-on experience of using WIA techniques was captured. Nonetheless, we used only one use case for both parts of the study. However, this was necessary to minimize participant variance, facilitate fair analyses comprehension, and ensure participants’ comfort in sharing their analyses without compromising sensitive information.
For the similar reasons, we used the marketing-specific use case even for non-marketing participants. To accommodate for this limitation, we provided examples of how the MMM task would be relevant to other departments, such as how MMM could enhance sales efficiency through optimized channel management and promotional strategies for sales managers. Additionally, we encouraged participants from the non-marketing background to draw parallels of the techniques to their own use cases, enriching our findings with other contextualized examples.
9. Conclusion
In this paper, we conducted a two-part study with 22 business users from real-world enterprises. The first interview study aimed to understand the WIA techniques employed, tools used, and challenges faced in making data-driven decisions. Our findings show that business users independently perform a range of advanced analyses due to various logistical constraints, but often rely on rudimentary methods that are insufficient for effective decision-making. To further learn business users’ feedback on benefits and opportunities of enhancing advanced analysis techniques, we conducted a follow-up task-based study with the same participants. In this study, they used four representative WIA techniques identified in the first interview study, that we implemented in a visual analytics prototype as a probe for them to use hands-on to make data-driven decisions. Our findings showed that these techniques improved decision-making speed and confidence while also highlighting the need for better dataset preparation, risk assessment, and domain knowledge integration. Finally, based on these insights, we make design recommendations for future business analytics systems.
Acknowledgements.
We thank Joe Hellerstein, Arjun Srinivasan, and the anonymous reviewers for their helpful feedback on the paper draft.